C784 Pre-Assessment Guide
Phillip December 9, 2025 No Comments

C784 Pre-Assessment Guide

C784 Pre-Assessment Guide

Name

 Western Governors University

C784 Applied Healthcare Statistics

Prof. Name

Date

1. Given the following set of numbers: 3.3, –2.3, –16, 25

What are the square roots of the whole number?

  • ± 2.3
  • ± 3.3
  • ± 5.0
  • ± 4.0

2. Given the following set of numbers: 10, 2.5, –4, –4.5, 5

What is the sum when adding together only the whole numbers?

  • 5.5
  • 11.0
  • 12.5
  • 15.0

3. Given the following expression: 4 + 7 × (6 – 3)²

What is the answer using proper order of operations?

  • 25
  • 67
  • 46
  • 99

4. Given the following expression: 5 × (30 – 27)² + 4

What is the correct answer using the proper order of operations?

  • 49
  • 34
  • 65
  • 81

5. What is the greatest common factor of 132 and the square root of 121 using prime factorization?

  • 1
  • 2
  • 10
  • 11

6. What is the greatest common factor of 24 and 84 using prime factorization?

  • 2
  • 4
  • 12
  • 8

9. A patient’s protein intake for today has been 15.76 g, 8.49 g, and 22.58 g.

What is the total protein intake rounded to the nearest tenth of a decimal?

  • 46.5 g
  • 46.8 g
  • 46.9 g
  • 47.0 g

10. Given the equation: (not fully shown)

What is the solution rounded to the nearest thousandth?

  • 0.03675
  • 0.0368
  • 0.368
  • 0.037

11. An intravenous drip is infused at a rate of 4.5 milliliters per minute.

How many milliliters are infused per hour?

  • 0.075 mL
  • 2.700 mL
  • 270.000 mL
  • 75.000 mL

12. A patient’s medication was increased by 40 milligrams.

What is the percent increase if the original medication was 180 milligrams?

  • 2.2%
  • 4.5%
  • 22.2%
  • 45.0%

13. What is the Fahrenheit equivalent for a patient’s room that needs to be kept at 25 degrees Celsius?

  • 70°
  • 77°
  • 75°
  • 80°

14. What is the Celsius equivalent for a patient who is running a fever of 102 degrees Fahrenheit?

  • 37.9°C
  • 38.9°C
  • 39.9°C
  • 40.9°C

15. What is the value of x in this equation? 2x + 3 = 13

  • 5
  • 10
  • 76
  • 106

24. A survey asked residents to rate their community hospital on a scale of 1 to 10, with 10 being the most happy with the hospital services. The data shows a positively skewed distribution. How are the mean, median and mode related?

  • Mode < Median < Mean
  • Mean < Median < Mode
  • Mean < Mode < Median
  • Median < Mode < Mean

25. An intensive care unit patient has frequent systolic blood pressure measures: 140, 150, 154, 154, 170, 176, 180

What is the mode?

  • 140.00
  • 154.00
  • 160.57
  • 170.00

26. Given data collected from research on cancer patients, what is the skew of the distribution?

  • Normal
  • Bimodal
  • Skewed right
  • Skewed left

27. Given this chart, what is true about the shape of the distribution?

  • The distribution is positively skewed.
  • The distribution is negatively skewed.
  • The distribution is symmetric.
  • The distribution is bimodal.

28. Given the pie chart with CCU, pediatric, surgery, and NICU, which percentage of patients acquired an infection in the surgery department?

  • 16%
  • 18%
  • 24%
  • 42%

29. The following weights were recorded for children: 93, 84, 70, 48, 59, 66, 88, 94, 84, 67, 74, 58, 82

Which method of display should be used to see the distribution and individual values?

  • Box plot
  • Stem plot
  • Bar chart
  • Pie chart

30. Heart rates recorded: 98, 88, 74, 83, 76, 99, 81, 73, 60, 66, 88, 94

Which display is appropriate to see exact values of individual observations?

  • Stem plot
  • Pie chart
  • Bar chart
  • Histogram

31. Hours of overtime worked by nurses: 3, 7, 2, 4

What should be used to display the data?

  • Box plot
  • Stem plot
  • Bar chart
  • Histogram

32. Given a pie chart showing percentages of hepatitis A, B, and C:

What causes this information to be misleading?

  • Hepatitis B should be bigger in size.
  • Hepatitis B should be smaller in size.
  • The percentages add up to more than 100%.
  • The percentages add up to less than 100%.

33. A clinic claims a dramatic increase in vaccines given over two years.

How is the clinic misrepresenting data?

  • By starting the vertical axis at 10
  • By showing only two years of data
  • By ending the vertical axis at 30
  • By creating a positive skew

C784 Pre-Assessment Guide


34. Which conclusion can be drawn from this distribution?

  • The distribution is positively skewed
  • The distribution is negatively skewed
  • The distribution is bimodal
  • The distribution is symmetric

35. A study on the relationship between gender and exercise type.

Which type of classification is this?

  • Categorical to categorical
  • Categorical to quantitative
  • Quantitative to quantitative
  • Quantitative to categorical

36. A study on the relationship between age and diastolic blood pressure.

Which type of classification is this?

  • Categorical to categorical
  • Quantitative to categorical
  • Categorical to quantitative
  • Quantitative to quantitative

37. A study on a categorical variable and a quantitative variable.

Which graphical display is appropriate?

  • Scatterplot
  • Two-way table
  • Decision table
  • Side-by-side box plot

38. Study on blood pressure among sleep apnea patients using CPAP and those who do not.

Which numerical measure is appropriate?

  • Regression analysis
  • Five-number summary
  • Correlation coefficient
  • Conditional percentages

39. Study on gender and hypertension.

Which numerical measure is appropriate?

  • Regression analysis
  • Five-number summary
  • Correlation coefficient
  • Conditional percentages

40. Study on blood pressure and hemoglobin A1c.

Which numerical measure is appropriate?

  • Standard deviation
  • Five-number summary
  • Correlation coefficient
  • Conditional percentages

41. Physician collects retrospective data on smoking and lung cancer.

Which data indicate the relationship?

  • Smokers are more likely than nonsmokers to have lung cancer because 77% is greater than 22%.
  • Nonsmokers are less likely because 78% is greater than 77%.
  • Nonsmokers are less likely because 23% is less than 77%.
  • Smokers are more likely because 23% is greater than 22%.

42. Study comparing surgery survival rates at two hospitals.

What is true about the relationship?

  • Hospital B has a higher survival rate because 98% is higher than 97%.
  • Hospital A has a higher survival rate because 3% is higher than 2%.
  • Hospital A has a higher survival rate because 97% is higher than 3%.
  • Hospital B has a higher survival rate because 98% is higher than 2%.

43. Study on physician attire preference by setting.

What is true about patient preference?

  • Patients prefer formal clothes and white coat because 67% > 17%.
  • Patients prefer scrubs because 65% > 17%.
  • Patients do not prefer scrubs because 7% < 9%.
  • Patients do not prefer scrubs instead of formal clothes because 65% < 67%.

44. Given a scatterplot:

What accurately describes the distribution and relationship?

  • Weak, no relationship
  • Strong, linear relationship
  • Weak, linear relationship
  • Strong, no relationship

45. Given a scatterplot:

What accurately describes the distribution and relationship?

  • Weak, linear relationship
  • Strong, linear relationship
  • Weak, nonlinear relationship
  • Strong, nonlinear relationship

46. Given a scatterplot with an outlier:

How does the outlier impact the relationship?

  • It weakens the negative relationship.
  • It weakens the positive relationship.
  • It strengthens the positive relationship.
  • It strengthens the negative relationship.

47. Given a scatterplot with an outlier:

What effect would removal of the outlier have?

  • It would strengthen a positive relationship.
  • It would weaken a positive relationship.
  • It would weaken a negative relationship.
  • It would strengthen a negative relationship.

48. Given a scatterplot:

What number of work days missed contains an outlier?

  • 3
  • 5
  • 8
  • 10

55. A pediatric practice has six physicians, four are male. A newborn patient is assigned randomly.

What is the probability assigned to a non-male physician?

  • 0.17
  • 0.33
  • 0.50
  • 0.67

56. Given scatterplot & regression equation, what is the anxiety score for someone practicing yoga 3 hours per day, 2 days per week?

  • 10
  • 8
  • 21
  • 26

57. Fair coin flipped again. Probability of heads?

  • 0.17
  • 0.2
  • 0.25
  • 0.5

58. Surveyed 100 patients; preferences: 35 Infusions, 30 Incision/drainage, 25 Wound care, 10 Vital signs.

Relative frequency for wound care?

  • 0.10
  • 0.25
  • 0.30
  • 0.35

59. Coin flipped 3 times, probability of exactly 2 heads?

  • 0.250
  • 0.375
  • 0.500
  • 0.875

60. Box with red, purple, blue balls. Two draws without replacement. Probability red ball is second?

  • 0.33
  • 0.50
  • 0.67
  • 0.83

61. Three pregnant women with one child each. Possible baby sex outcomes (M or F)?

  • MMM, MMF, MFM, MFF, FMM, FMF, FFM, FFF

62. Patient has elevated temperature 20%, high cholesterol 72%, both 14.4%. What’s true?

  • The values are independent (0.2 × 0.72 = 0.144)

63. Cold climates: 30% get influenza, 15% appendicitis. Probability of both (approximate)?

  • 5%
  • 15%
  • 30%
  • 45%

64. Probability of running outside 0.45, joining gym 0.40, both 0.25. Probability either running or gym?

  • 0.60
  • 0.65
  • 0.70
  • 0.85

65. 27 patients: 7 have diabetes, 9 have heart disease, 3 both.

How to find probability of diabetes or heart disease?

  • P(D or H) = P(D) + P(H) − P(D and H) = (7/27) + (9/27) − (3/27)

66. Clinic sees 8 patients/hour; probability patient has high BP = 0.63; probability none have high BP?

  • (1 − 0.63)^8 = (0.37)^8

67. 625 group physician practices: 300 primary care, 225 specialty care, 100 both.

Probability of primary or specialty?

  • (300/625) + (225/625) − (100/625)

68. 35% patients Medicare, 15% Medicaid; 10% of Medicare patients also have Medicaid.

Probability patient has both?

  • 0.35 × 0.10 = 0.035

69. 14% get Virus A, 18% Virus B, 10.62% both.

Probability that a person who contracts Virus B also contracts Virus A?

  • P(A|B) = 0.1062 / 0.18 ≈ 0.59 or 59%

70. 37% patients had sore throat, 36% itchy eyes, 51% of itchy eyes patients had sore throat.

Probability patient had both?

  • 0.36 × 0.51 = 0.1836 or 18.36%

 

error: Content is protected, Contact team if you want Free paper for your class!!